4. A current of 4 A flows through a resistor, and the voltage across the resistor is 32 V.
a) What is the power dissipated by the resistor?
b) How much energy is dissipated by the resistor in 20 s?
c) What is the resistance of the resistor?
1
Expert's answer
2020-04-10T08:41:14-0400

(a) the power dissipated by the resistor

"P=IV=(4\:\rm A)\times (32\:\rm V)=128\:\rm W."

(b) the amount of energy dissipated by the resistor in 20 s

"E=Pt=(128\:\rm W)\times (20\:\rm s)=2560\:\rm J."

(c) the resistance of the resistor

"R=\frac{V^2}{P}=\frac{(32\:\rm V)^2}{(128\:\rm W)}=8\:\rm \Omega."